OpenStudy (anonymous):

A single die is rolled twice. Find the probability of rolling a 2 the first time and a 3 the second time.

OpenStudy (kropot72):

There are 6 numbers on a die. The probability of rolling any of the number is 1/6. Therefore the probability of rolling a 2 the first time and a 3 the second time is given by: \[\large P(2\ then\ 3)=\frac{1}{6}\times\frac{1}{6}=you\ can\ calculate\]

OpenStudy (kropot72):

numbers*